OpenStudy (anonymous):

10. Solve the proportion. 5/8 = y/40 A.y = 25 B.y = 13 C.y = 11 D.y = 10 @Destinymasha almost done (:

OpenStudy (destinymasha):

25

OpenStudy (destinymasha):

8 x 5 = 40, and 5x5 = 25

jimthompson5910 (jim_thompson5910):

5/8 = y/40 5*40 = 8*y ... cross multiply 200 = 8y 8y = 200 y = 200/8 y = 25